Nnesha (nnesha):

o^_^o ello! trig trig trig trig Rewrite as an equivalent expression that does not contain powers of trig functions greater than 1. \[\huge\rm 8 \cos^2 x\]

OpenStudy (anonymous):

you mean to tell me you dont know how to do this

OpenStudy (anonymous):

You was teaching me conics

Nnesha (nnesha):

;-;

OpenStudy (anonymous):

2 cos² θ = cos 2θ + 1

OpenStudy (anonymous):

cos 2θ = cos² θ - sin² θ

Nnesha (nnesha):

cos^2 theta = 1- 2 sin^2 theta

Nnesha (nnesha):

okay so answer is 4+4cos2x but how ???

OpenStudy (anonymous):

:)

OpenStudy (rational):

\[8\cos^2x = 4\cdot 2\cos^2x = 4\cdot (1+\cos (2x)) = 4+4\cos (2x)\]

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

I didn't want to give a sort of direct answer

Nnesha (nnesha):

ohhh \[\huge\rm cos (2x) = 2 \cos^2 x -1\] solve for 2cos^2 \[\rm cos (2x) +1 = 2\cos^2\] hmmm that's how 2cos^2 = cos (2x) +1

OpenStudy (anonymous):

yup :)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @DaWizjr I didn't want to give a sort of direct answer \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @Nnesha okay so answer is 4+4cos2x but how ??? \(\color{blue}{\text{End of Quote}}\) :D i already know the answer :-)

Nnesha (nnesha):

thanks!!! rational & Dawizjr!!!

OpenStudy (paki):

@Nnesha mujhay math nahi ati :(